OpenStudy (anonymous):

solve the system of equations using any method? x+y=9 x-3y=-3

OpenStudy (anonymous):

Which method do you want to use?

OpenStudy (lgbasallote):

i would use elimination if i were you

OpenStudy (anonymous):

Hey don't give in to peer pressure. You pick whatever you want.

OpenStudy (anonymous):

it doesnt matter its up too you!!

OpenStudy (lgbasallote):

but peer pressure is fun!!!

OpenStudy (anonymous):

@Spacelimbus ??

OpenStudy (anonymous):

How may I help? (-: I believe there are only two methods that can be put into consideration at this problem. One is substitution, meaning you express one variable in respect of the other and/or elimination, which means that you're trying to 'eliminate' one variable. I, personally, would use substitution (-: Fortunately there is no right or wrong here.

OpenStudy (anonymous):

@RyanL. ?

OpenStudy (anonymous):

need an answer @Callisto

OpenStudy (callisto):

Subtract the second equation from the first equation, what do you get?

OpenStudy (anonymous):

idk??

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

??

OpenStudy (callisto):

(x+y) - (x-3y) =9 - (-3) <- Can you solve it?

OpenStudy (anonymous):

nooo

OpenStudy (callisto):

Consider the left side, (x+y) - (x-3y) = x + y - x + 3y = ...?

OpenStudy (callisto):

Blah.... Is the question 'which method would you use to solve the equations' or 'solve the questions'?

OpenStudy (anonymous):

solve the system of equations using any method!!

OpenStudy (jiteshmeghwal9):